### v1

@@ Appendix C Upper Bound of Randomized Response Method
By Lemma 11, with probability at least $41-\beta/$ 4, we have that each of the absolute values $\left|Y_{(i,j)}\right|$’s is at most $L=O(1/\varepsilon\cdot\log(n/\beta))$. We then bound the three terms above conditioning on this event happening:
The first term $w_{\max}^{2}\sum_{(i,j,k)\in\binom{V}{3}}q_{(i,j,k)}(Y_{(i,j)}+Y_{(j,k)}+Y_{(k,i)})$ in Equation 28 can be bounded by using a Chernoff bound. Specifically, since $\sum_{(i,j,k)\in\binom{V}{3}}q_{(i,j,k)}(Y_{(i,j)}+Y_{(j,k)}+Y_{(k,i)})=\sum_{(i,j)\in\binom{V}{2}}\sum_{k\in V}q_{(i,j,k)}Y_{(i,j)}$, it holds that $\sum_{(i,j)\in\binom{V}{2}}\sum_{k\in V}q_{(i,j,k)}Y_{(i,j)}$ is the sum of $\binom{n}{2}$-many independent random variables with mean 00 and bounded by $[-nL,nL]$. Thus, by Chernoff bound,
@@
We first consider the second term in Equation 28. The martingale considered is as follows: Given an arbitrary order of the $\binom{n}{2}$ pairs, namely, $e_{1},\dots,e_{\binom{n}{2}}$. For the sake of simplicity, we denote $\mathcal{E}_{i}$ as $\{e_{1},e_{2},\dots,e_{i}\}$. Then we let $X_{0}=0$, and let
@@
For the third term in Equation 28, we can define the martingale as:
@@
for any triangle-motif cut query with probability at least $41-\beta/$ 4.

### v2

@@ Appendix C Upper Bound of Randomized Response Method
By Lemma 11, with probability at least $1-\beta/4$, we have that each of the absolute values $\left|Y_{(i,j)}\right|$’s is at most $L=O(1/\varepsilon\cdot\log(n/\beta))$. We then bound the three terms above conditioning on this event happening:
The first term $w_{\max}^{2}\sum_{(i,j,k)\in\binom{V}{3}}q_{(i,j,k)}(Y_{(i,j)}+Y_{(j,k)}+Y_{(k,i)})$ in Equation˜20 can be bounded by using a Chernoff bound. Specifically, since $\sum_{(i,j,k)\in\binom{V}{3}}q_{(i,j,k)}(Y_{(i,j)}+Y_{(j,k)}+Y_{(k,i)})=\sum_{(i,j)\in\binom{V}{2}}\sum_{k\in V}q_{(i,j,k)}Y_{(i,j)}$, it holds that $\sum_{(i,j)\in\binom{V}{2}}\sum_{k\in V}q_{(i,j,k)}Y_{(i,j)}$ is the sum of $\binom{n}{2}-$many independent random variables with mean 0 and bounded by [nL,nL][-nL,nL]. Thus, by Chernoff bound,
@@
We first consider the second term in Equation˜20. The martingale considered is as follows: Given an arbitrary order of the $\binom{n}{2}$ pairs, namely, $e_{1},\dots,e_{\binom{n}{2}}$. For the sake of simplicity, we denote $\mathcal{E}_{i}$ as $\{e_{1},e_{2},\dots,e_{i}\}$. Then we let $X_{0}=0$, and let
@@
$\mathbb{E}\left[X_{i}|X_{0},X_{1},\dots,X_{i-1}\right]$
$=$ $\mathbb{E}\left[X_{i}|e_{1},e_{2},\dots,e_{i-1}\right]$
$=$ $X_{i-1}+\mathbb{E}\left[Y_{e_{i}}|Y_{e_{1}},Y_{e_{2}},\dots,Y_{e_{i-1}}\right]\cdot\left(\sum_{(i_{1},i_{3})\in\mathcal{E}_{i-1}}q_{(i_{1},i_{2},i_{3})}Y_{(i_{1},i_{3})}+\sum_{(i_{2},i_{3})\in\mathcal{E}_{i-1}}q_{(i_{1},i_{2},i_{3})}Y_{(i_{2},i_{3})}\right)$
$=$ $X_{i-1}.$
@@
For the third term in Equation˜20, we can define the martingale as:
@@
for any triangle-motif cut query with probability at least $1-\beta/4$.